# Neural ordinary differential equations for model order reduction of stiff systems

### v1

@@ 4.5 Test case 5: POLLU problem
where the reactions $\boldsymbol{r}$ are connected to $\boldsymbol{u}$ and $\boldsymbol{\mu}^{0}$ by rates reported in Table 5. The initial condition is
@@
Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{0}$ Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{0}$ Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{0}$
@@
Table 5: Reference reaction rates for the POLLU model that define $\boldsymbol{\mu}^{0}$.
It is usually considered that the time interval $t\in[0,60)$ is representative of the behavior of the reactants sufficiently. The reference solution is obtained by using a tolerance of $10^{-10}$. Full details about the model can be found in [63]. The parameter space $\Gamma$ is built by changing the three components of $\boldsymbol{\mu}$ that most impact the system. Namely, a preliminary sensibility analysis has shown that the components $4,6$, and $14$ of $\boldsymbol{\mu}$ are the most relevant, that is they produced the largest variation in the solution $\boldsymbol{u}$ when perturbed. Hence, we define the parameter space as
$\Gamma=\prod_{i\notin\{4,6,14\}}\{\mu_{i}^{0}\}\times\prod_{i\in\{4,6,14\}}\left[\frac{1}{2}\mu_{i}^{0},2\mu_{i}^{0}\right],$
where $\boldsymbol{\mu}^{0}$ is the reference parameter value used for the system, reported in Table 5. The intervals are discretized with 16 uniform points in each coordinate direction. The validation dataset is made by the midpoints of the training datasets.
The model is tested for two parameters $\boldsymbol{\mu}^{1}_{test},\boldsymbol{\mu}^{2}_{test}$ which are equal to $\boldsymbol{\mu}^{0}$ apart from the three most relevant components, namely $\mu_{test,i}^{1}=0.525\mu_{i}^{0},\,\mu_{test,i}^{2}=1.975\mu_{i}^{0},\,i=4,6,14$. Results are shown in Figure 10. The ROM is able to follow the reference solution across all the reactions. A quantitative comparison with the Radau solver is reported in Table 6. The number of right-hand side evaluations of the explicit solver is similar to the implicit one. On the other hand, each evaluation of the right-hand side, each evaluation of the Jacobian, and each LU decomposition is more costly than in the previous test cases since $N_{u}$ is larger. However, the cost of evaluating the neural network is similar to the previous test cases since the dimensions of the networks are comparable. Indeed, the cost of the ROM is still smaller than the cost of the Radau solver while keeping comparable accuracy. These results are of particular significance because they prove that as the state space gets larger our methodology becomes more cost effective.
@@ 5 Conclusions
In this work, we have developed a novel methodology to tackle the stiffness in neural ODEs. The approach hinges on a suitable time reparametrization of the system that significantly reduces the stiffness, making it possible to efficiently apply explicit solvers to the neural ODEs. The construction of the time reparametrization is general and completely data-driven, making it ideal for an application to ROMs.
By leveraging the intrinsic capabilities of neural ODEs to handle continuous-time data, we capture the dynamics of stiff systems effectively. The accuracy, robustness, and efficiency of the methodology were tested in five famous benchmark test cases from the literature. Namely, when applied to periodic systems, the model generalizes to data well beyond the time training interval. We attribute this characteristic primarily to the autonomous nature of the neural ODE and the time reparametrization model. When applied to chemical equations, our model always maintained the positivity of the concentration, even when applied to problems suffering from numerical instability. Moreover, the work/precision tradeoff of our models always proved comparable or favorable when confronted with a state-of-the-art Runge-Kutta implicit solver of the Radau II A kind.
In conclusion, our work highlights the promising potential of neural ODEs in creating efficient and accurate reduced-order models for stiff ODEs. Future research is directed towards improving the accuracy of the time mapping that, as of right now, stands as the major accuracy bottleneck of the ROM. We also aim to test this framework to handle the semi-discrete formulation of partial differential equations and other higher-dimensional systems. Moreover, techniques like model distillation, network pruning, and better hyperparameter tuning could further reduce the computational cost of our approach.

### v2

@@ 4.5 Test case 5: POLLU problem
where the reactions $\boldsymbol{r}$ are connected to $\boldsymbol{u}$ and $\boldsymbol{\mu}^{(0)}$ by rates reported in Table 5. The initial condition is
@@
Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{(0)}$ Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{(0)}$ Reaction $\boldsymbol{r}$ $\boldsymbol{\mu}^{(0)}$
@@
Table 5: Reference reaction rates for the POLLU model that define $\boldsymbol{\mu}^{(0)}$.
It is usually considered that the time interval $t\in[0,60]$ is representative of the behavior of the reactants sufficiently. The reference solution is obtained by using a tolerance of $10^{-10}$. Full details about the model can be found in [73]. The parameter space $\Gamma$ is built by changing the three components of $\boldsymbol{\mu}$ that most impact the system. Namely, a preliminary sensibility analysis has shown that the components $4,6$, and $14$ of $\boldsymbol{\mu}$ are the most relevant, that is they produced the largest variation in the solution $\boldsymbol{u}$ when perturbed. Hence, we define the parameter space as
$\Gamma=\bigtimes_{i\notin\{4,6,14\}}\{\mu_{i}^{(0)}\}\times\bigtimes_{i\in\{4,6,14\}}\left[\frac{1}{2}\mu_{i}^{(0)},2\mu_{i}^{(0)}\right],$
where $\boldsymbol{\mu}^{(0)}$ is the reference parameter value used for the system, reported in Table 5. The intervals are discretized with 16 uniform points in each coordinate direction. The validation dataset is made by the midpoints of the training datasets.
The model is tested for two parameters $\boldsymbol{\mu}^{\text{test}(1)},\boldsymbol{\mu}^{\text{test}(2)}$ which are equal to $\boldsymbol{\mu}^{(0)}$ apart from the three most relevant components, namely $\mu^{\text{test}(1)}_{i}=0.525\mu_{i}^{(0)},\,\mu_{i}^{\text{test}(2)}=1.975\mu_{i}^{(0)},\,i=4,6,14$. Results are shown in Figure 13. The ROM is able to follow the reference solution across all the reactions. A quantitative comparison with the Radau solver is reported in Table 6. The number of right-hand side evaluations of the explicit solver is similar to the implicit one. On the other hand, each evaluation of the right-hand side, each evaluation of the Jacobian, and each LU decomposition is more costly than in the previous test cases since $N_{u}$ is larger. However, the cost of evaluating the neural network is similar to the previous test cases since the dimensions of the networks are comparable. Indeed, the cost of the ROM is still smaller than the cost of the Radau solver while keeping comparable accuracy. These results are of particular significance because they prove that as the state space gets larger our methodology becomes more cost effective.
@@ 5 Conclusions
In this work, we have developed a novel methodology to tackle some kinds of stiffness in neural ODEs. The approach hinges on a suitable time reparametrization of the system that significantly reduces the stiffness, making it possible to efficiently apply explicit solvers to the neural ODEs. The construction of the time reparametrization is general and completely data-driven, making it ideal for an application to ROMs.
By leveraging the intrinsic capabilities of neural ODEs to handle continuous-time data, we capture the dynamics of stiff systems effectively. The accuracy, robustness, and efficiency of the methodology were tested in five famous benchmark test cases from the literature. Namely, when applied to periodic systems, the model generalizes to data well beyond the time training interval. We attribute this characteristic primarily to the autonomous nature of the neural ODE and the time reparametrization model. When applied to chemical equations, our model always maintained the positivity of the concentration, even when applied to problems suffering from numerical instability. Moreover, the work/precision tradeoff of our models often proved comparable or favorable when confronted with a state-of-the-art Runge-Kutta implicit solver of the Radau II A kind. In a few cases where this was not observed, our model provided a computationally cheaper alternative for approximate solutions while preserving concentration positivity, which is not possible with the Radau solver for small tolerances.
In conclusion, our work highlights the promising potential of neural ODEs in creating efficient and accurate reduced-order models for stiff ODEs. Stiffness characterized by reasonably large eigenvalues but very high frequency remains a challenge for our approach, and it is an active topic of research. Future research is directed towards improving the accuracy of the time mapping that, as of right now, stands as the major accuracy bottleneck of the ROM. We also aim to test this framework to handle the semi-discrete formulation of partial differential equations and other higher-dimensional systems. Moreover, techniques like model distillation, network pruning, and better hyperparameter tuning could further reduce the computational cost of our approach.